Kareena obtained a $500,000 housing loan. The loan interest rate is 6% per annum and the agreed loan repayment period is 25 years. Calculate Kareena’s monthly mortgage payment.
Answer #1

We will use Excel's PMT function to arrive at monthly mortgage payment

Loan Amount = $500,000

Interest Rate = 6% p.a

Monthly Interest Rate = (6/12)% = 0.5%

Repayment period = 25 years

Total no of months = (25 * 12) = 300 months

A B C Loan Amount Monthly Interest Rate No of periods Monthly Payment $500,000 0.50% 300 ($3,221.51)

А в Loan Amount 500000 Monthly Interest Rate =6%/12 No of periods =25*12 Monthly Payment =PMT(C4,C5,C3,0)

Monthly Mortgage payment = $3,221.51